# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
Among the existing hyperbolic models, we chose to work with the Poincaré half-space model because (i) it is natural to represent diffusion times on a dyadic grid in this model, following previous work [Lin et al., 2023], and (ii) there exists a closed-form solution of the hyperbolic LCA and the HD-LCA, which are the key components in our tree decoding algorithm in Alg. 1. Since the four standard models of hyperbolic space [Bowditch, 2007]: the Poincaré disk, Lorentz, Poincaré half-space, and Beltrami-Klein, are equivalent and isometric, it is possible to use the closed-form solution of LCA in the alternative models, e.g., the Lorentz model for numerical stability. However, we did not encounter any numerical instabilities in our empirical studies using our implementation in the Poincaré half-space model.
@@
We consider hyperbolic space with negative constant curvature of $-1$ because it aligns with the common practice in hyperbolic representation [Nickel and Kiela, 2017, 2018, Sala et al., 2018, Chami et al., 2020]. Any negative constant curvature can indeed be used and considered as a hyperparameter, e.g., [Chami et al., 2019]. The modifications that it will require from our algorithm are minimal, and most of the steps can be applied seamlessly in such a space. We conducted experiments on scRNA-seq data to investigate the role of curvature in our method. Specifically, we vary the curvature and explore the effects of constant negative curvature $-1/\kappa$ where $\kappa>0$. Fig. 7 presents the classification accuracy of the Zeisel and CBMC datasets as a function of the curvature. We see that decreasing the curvature slightly improves the cell classification accuracy. This suggests that while the choice of curvature does have a large impact, our method remains robust across a range of curvatures. In addition, fine-tuning the curvature can enhance performance, and starting with the default choice of $-1/\kappa=-1$ is a reasonable and commonly accepted practice.
@@
We note that other TWD baselines could use other two-dimensional Poincaré representations (e.g., Poincaré embedding (PE) [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019]) and decode the hyperbolic embedding into a tree by HHC [Chami et al., 2020]. We denoted these baselines by PE-HHC-TWD and HVAE-HHC-TWD, respectively, and present their results in Tab. E.3.5. We see in the table that HHC [Chami et al., 2020] applied to a PE [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019] is less effective than our TWD by a large margin. This discrepancy is attributed to the sensitivity of the recovery of the hidden feature hierarchy using the competing methods, a finding that is consistent with previous works [Sala et al., 2018, Sonthalia and Gilbert, 2020, Lin et al., 2023].
@@ E.3.6 Comparison to Non-TWD Baselines
For non-TWD baselines, we also compare our method to Word Mover’s Distance (WMD) [Kusner et al., 2015] for word-document datasets and Gene Mover’s Distance (GMD) [Bellazzi et al., 2021] for scRNA-seq datasets in Sec. 5. Note that WMD [Kusner et al., 2015] computes the Wasserstein distance using the Euclidean distance $\widetilde{d}$ between precomputed Word2Vec embeddings [Mikolov et al., 2013] as the ground cost, while GMD [Bellazzi et al., 2021] is defined similarly, using the Euclidean distance $\widetilde{d}$ between precomputed Gene2Vec embeddings as the ground cost. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies [Kusner et al., 2015, Huang et al., 2016, Le et al., 2019, Takezawa et al., 2021, Yamada et al., 2022, Chen et al., 2024, Bellazzi et al., 2021].
@@
It is important to note that using the ground metric $\widetilde{d}$ (e.g., Euclidean) in WMD, GMD, or existing TWD methods results in the (approximated) Wasserstein distance being based on the predefined metric $\widetilde{d}$. In contrast, our method employs an initial distance $d$, computed using the cosine similarity in the experiments, to compute an affinity matrix within the diffusion operator, which is central to kernel and manifold learning methods. This process leads to a Wasserstein distance based on a latent hierarchical metric induced by $d$. As noted in App. D.1, the kernel type, scale parameter, and initial distance metric selection in our approach significantly influence its effectiveness. These choices are task-specific and are critical in achieving optimal performance [Lindenbaum et al., 2020a].
@@
In addition, we conduct additional tests evaluating classification performance directly in the feature space, without using Wasserstein distance or TWD. Tab. E.3.6 presents the document and cell classification accuracy in the feature space using the Euclidean distance and cosine distance. We see that the classification performance relying on generic metrics (Euclidean or cosine) is much less effective than WMD and GMD, which aligns with the findings reported in WMD [Kusner et al., 2015] and GMD [Bellazzi et al., 2021]. This highlights that incorporating the feature relationship into a distance between samples improves the effectiveness of distance metric learning. Furthermore, we see that our TWD method performs better than the distance relying on generic metrics, WMD and GMD. This suggests that inferring explicitly the feature hierarchies is important and leads to the effectiveness of distance metric learning for data with a hierarchical structure.
@@ F.1 Single-scale Hyperbolic Diffusion Binary Tree Decoding
In Alg. 1, the tree decoding is based on the multi-scale embeddings and the corresponding HD-LCAs. In Alg. 3 we present a similar algorithm based on the single scale hyperbolic embedding $\mathbf{z}_{j}^{k}$ and the corresponding $k$-th hyperbolic LCAs ${\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}}$ in Def. 4.1, giving rise to a (single-scale) binary tree $B^{k}$. Specifically, we merge the pair of features exhibiting the highest similarity at the $k$-th hyperbolic LCA, determined by the smallest value of $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})$ in Prop. 4.1. Following Def. 4.2, only one of the hyperbolic LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathbb{H}^{m+1}}^{k}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathbb{H}^{m+1}}^{k}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathbb{H}^{m+1}}^{k}$, holds in $B^{k}$. The merged nodes are then connected by edges whose weights are assigned by $d_{\mathbb{H}^{m+1}}(\cdot,\cdot)$ between the corresponding points. It is worth noting that considering $(K_{c}+1)$ trees $\{B^{k}\}_{k=0}^{K_{c}}$ generated by Alg. 3 gives rise to a tree-sliced Wasserstein distance as defined in Prop. 4.5. In Prop. 4.5, we show that our TWD efficiently captures the tree-sliced information encoded within these multi-scale trees, circumventing the need for their explicit construction. In addition, we posit that our high-dimensional tree decoding in Alg. 3 for single-scale hyperbolic embedding can be readily applied to other hyperbolic embeddings, e.g., [Nickel and Kiela, 2017, 2018, Nagano et al., 2019], extending its applicability and utility. While this extension is technically feasible, it does not carry the theoretical guarantees as those established in Thm. 4.1 and Thm. 4.2.
@@ F.2 Putting the Features on the Leaves of the Tree
The ground truth latent tree $T$ does not assume that features are positioned at the leaves. In addition, note that the hyperbolic embedding [Lin et al., 2023], which is consistent with other common hyperbolic embedding methods [Chamberlain et al., 2017, Nickel and Kiela, 2017, 2018, Sala et al., 2018], does not restrict all the features to the leaves. Assigning the features to the leaf nodes is only carried out in the tree decoding (Alg. 1), which is in line with other tree decoding algorithms [Sarkar, 2011, Chami et al., 2020, Indyk and Thaper, 2003]. Note that our tree decoding algorithm generates a weighted binary tree. We assert that this does not pose a restriction because any tree can be effectively transformed into a binary tree [Bowditch, 2007, Cormen et al., 2022]. For instance, consider the flexible tree in Fig. 10 (a): [sport[soccer,football,basketball]] (where “sports” is a parent node with three children: “basketball”, “soccer”, and “football”). This tree can be accurately modeled as a (weighted) binary tree as in Fig. 10 (b): [$c$[$b$[$a$[soccer, football], basketball]], sport] (where “sports”, “basketball”, “soccer”, and “football” are all leaves, $a,b,c$ are internal nodes, and the edges encode the same relationships as the flexible tree above).
@@ F.3 Differentiability of Our TWD
The proposed TWD can be expressed in a matrix form [Takezawa et al., 2021] as
@@
Note that the hyperbolic embedding in Lin et al. [2023] can be recast as a function mapping from the feature diffusion operator $\mathbf{P}$ to the product manifold of the hyperbolic spaces $\mathcal{M}$, denoted by $f_{1}:\mathbf{P}\rightarrow\mathcal{M}$. The weight $\alpha_{v}$ in the our TWD can then be expressed by
@@
F.4 The Differences between Our TWD and the Embedding Distance [Lin et al., 2023]
Our TWD is a TWD between samples that incorporates the hidden hierarchical structure of the features, while the embedding distance [Lin et al., 2023] is a hierarchical distance between the samples. Namely, in our TWD, the features are assumed to have hidden hierarchies, whereas in Lin et al. [2023], the samples are assumed to have hidden hierarchies. The work in Lin et al. [2023] does not involve TWD or OT.
More concretely, in comparison to the work in Lin et al. [2023], our work presents a new algorithm for tree decoding (Alg. 1) and a corresponding new TWD (Alg. 2). Our work also includes theoretical analysis of these two algorithms:
@@
Seemingly, from a metric learning viewpoint, both our TWD and the embedding distance [Lin et al., 2023] generate distances between the samples. Since the two are fundamentally different, the comparison can be carried out in two ways: (i) directly compute the embedding distance between the samples (denoted as ED), and (ii) use the embedding distance between the features as the ground distance for the computation of the Wasserstein distance between the samples (denoted as OT-ED). We present the classification accuracy of the word-document and scRNA-seq datasets of these two baselines in Tab. F.4. We see a clear empirical advantage of our TWD in classifying documents and scRNA-seq data compared to embedding distance (ED), as our TWD learns the latent hierarchy of the features explicitly by constructing a tree, and then, incorporating it in a meaningful distance between the samples. We note that while OT-ED performs similarly to our TWD in terms of accuracy (empirically corroborating Thm. 4.2), the computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our proposed TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. This computational advantage is empirically validated and illustrated in Fig. 4.
@@ F.5 Intuition Behind Single-Linkage Tree Construction in Hyperbolic Embeddings
The single-linkage tree aligns well with hyperbolic embeddings because it inherently emphasizes hierarchical relationships by connecting points based on the measure of similarity determined by their HD-LCAs, where there is the “treelike” nature of a hyperbolic space [Bowditch, 2007]. Specifically, Prop. 4.4 shows that the LCA depth closely approximates the Gromov product, which corresponds to the tree depth in 0-hyperbolic metrics.
In the context of tree-Wasserstein distance (TWD), many tree construction methods aim to approximate a given ground metric $\widetilde{d}$ (often Euclidean) using tree metrics, enabling efficient computation of optimal transport distances. The quality of TWD methods is typically assessed by the mean relative error between the proposed TWD and the Wasserstein distance with the ground metric $\widetilde{d}$. For instance, an ultrametric incurs a $\log m$ distortion with respect to Euclidean distance [Fakcharoenphol et al., 2003], as utilized in UltraTree [Chen et al., 2024]. Hence, the ultrametric Wasserstein distance [Chen et al., 2024] incurs a $\log m$ distortion compared to the original Euclidean Wasserstein distance.

### v3

@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
Among the existing hyperbolic models, we chose to work with the Poincaré half-space model because (i) it is natural to represent diffusion times on a dyadic grid in this model, following previous work (Lin et al., 2023), and (ii) there exists a closed-form solution of the hyperbolic LCA and the HD-LCA, which are the key components in our tree decoding algorithm in Alg. 1. Since the four standard models of hyperbolic space (Bowditch, 2007): the Poincaré disk, Lorentz, Poincaré half-space, and Beltrami-Klein, are equivalent and isometric, it is possible to use the closed-form solution of LCA in the alternative models, e.g., the Lorentz model for numerical stability. However, we did not encounter any numerical instabilities in our empirical studies using our implementation in the Poincaré half-space model.
@@
We consider hyperbolic space with negative constant curvature of $-1$ because it aligns with the common practice in hyperbolic representation (Nickel & Kiela, 2017; 2018; Sala et al., 2018; Chami et al., 2020). Any negative constant curvature can indeed be used and considered as a hyperparameter, e.g., (Chami et al., 2019). The modifications that it will require from our algorithm are minimal, and most of the steps can be applied seamlessly in such a space. We conducted experiments on scRNA-seq data to investigate the role of curvature in our method. Specifically, we vary the curvature and explore the effects of constant negative curvature $-1/\kappa$ where $\kappa>0$. Fig. 7 presents the classification accuracy of the Zeisel and CBMC datasets as a function of the curvature. We see that decreasing the curvature slightly improves the cell classification accuracy. This suggests that while the choice of curvature does have a large impact, our method remains robust across a range of curvatures. In addition, fine-tuning the curvature can enhance performance, and starting with the default choice of $-1/\kappa=-1$ is a reasonable and commonly accepted practice.
@@
We note that other TWD baselines could use other two-dimensional Poincaré representations (e.g., Poincaré embedding (PE) (Nickel & Kiela, 2017) or hyperbolic VAE (Nagano et al., 2019)) and decode the hyperbolic embedding into a tree by HHC (Chami et al., 2020). We denoted these baselines by PE-HHC-TWD and HVAE-HHC-TWD, respectively, and present their results in Tab. E.3.5. We see in the table that HHC (Chami et al., 2020) applied to a PE (Nickel & Kiela, 2017) or hyperbolic VAE (Nagano et al., 2019) is less effective than our TWD by a large margin. This discrepancy is attributed to the sensitivity of the recovery of the hidden feature hierarchy using the competing methods, a finding that is consistent with previous works (Sala et al., 2018; Sonthalia & Gilbert, 2020; Lin et al., 2023).
@@ E.3.6 Comparison to Non-TWD Baselines
For non-TWD baselines, we also compare our method to Word Mover’s Distance (WMD) (Kusner et al., 2015) for word-document datasets and Gene Mover’s Distance (GMD) (Bellazzi et al., 2021) for scRNA-seq datasets in Sec. 5. Note that WMD (Kusner et al., 2015) computes the Wasserstein distance using the Euclidean distance $\widetilde{d}$ between precomputed Word2Vec embeddings (Mikolov et al., 2013) as the ground cost, while GMD (Bellazzi et al., 2021) is defined similarly, using the Euclidean distance $\widetilde{d}$ between precomputed Gene2Vec embeddings as the ground cost. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies (Kusner et al., 2015; Huang et al., 2016; Le et al., 2019; Takezawa et al., 2021; Yamada et al., 2022; Chen et al., 2024; Bellazzi et al., 2021).
@@
It is important to note that using the ground metric $\widetilde{d}$ (e.g., Euclidean) in WMD, GMD, or existing TWD methods results in the (approximated) Wasserstein distance being based on the predefined metric $\widetilde{d}$. In contrast, our method employs an initial distance $d$, computed using the cosine similarity in the experiments, to compute an affinity matrix within the diffusion operator, which is central to kernel and manifold learning methods. This process leads to a Wasserstein distance based on a latent hierarchical metric induced by $d$. As noted in App. D.1, the kernel type, scale parameter, and initial distance metric selection in our approach significantly influence its effectiveness. These choices are task-specific and are critical in achieving optimal performance (Lindenbaum et al., 2020a).
@@
In addition, we conduct additional tests evaluating classification performance directly in the feature space, without using Wasserstein distance or TWD. Tab. E.3.6 presents the document and cell classification accuracy in the feature space using the Euclidean distance and cosine distance. We see that the classification performance relying on generic metrics (Euclidean or cosine) is much less effective than WMD and GMD, which aligns with the findings reported in WMD (Kusner et al., 2015) and GMD (Bellazzi et al., 2021). This highlights that incorporating the feature relationship into a distance between samples improves the effectiveness of distance metric learning. Furthermore, we see that our TWD method performs better than the distance relying on generic metrics, WMD and GMD. This suggests that inferring explicitly the feature hierarchies is important and leads to the effectiveness of distance metric learning for data with a hierarchical structure.
@@ F.1 Single-scale Hyperbolic Diffusion Binary Tree Decoding
In Alg. 1, the tree decoding is based on the multi-scale embeddings and the corresponding HD-LCAs. In Alg. 3 we present a similar algorithm based on the single scale hyperbolic embedding $\mathbf{z}_{j}^{k}$ and the corresponding $k$-th hyperbolic LCAs ${\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}}$ in Def. 4.1, giving rise to a (single-scale) binary tree $B^{k}$. Specifically, we merge the pair of features exhibiting the highest similarity at the $k$-th hyperbolic LCA, determined by the smallest value of $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})$ in Prop. 4.1. Following Def. 4.2, only one of the hyperbolic LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathbb{H}^{m+1}}^{k}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathbb{H}^{m+1}}^{k}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathbb{H}^{m+1}}^{k}$, holds in $B^{k}$. The merged nodes are then connected by edges whose weights are assigned by $d_{\mathbb{H}^{m+1}}(\cdot,\cdot)$ between the corresponding points. It is worth noting that considering $(K_{c}+1)$ trees $\{B^{k}\}_{k=0}^{K_{c}}$ generated by Alg. 3 gives rise to a tree-sliced Wasserstein distance as defined in Prop. 4.5. In Prop. 4.5, we show that our TWD efficiently captures the tree-sliced information encoded within these multi-scale trees, circumventing the need for their explicit construction. In addition, we posit that our high-dimensional tree decoding in Alg. 3 for single-scale hyperbolic embedding can be readily applied to other hyperbolic embeddings, e.g., (Nickel & Kiela, 2017; 2018; Nagano et al., 2019), extending its applicability and utility. While this extension is technically feasible, it does not carry the theoretical guarantees as those established in Thm. 4.1 and Thm. 4.2.
@@ F.2 Putting the Features on the Leaves of the Tree
The ground truth latent tree $T$ does not assume that features are positioned at the leaves. In addition, note that the hyperbolic embedding (Lin et al., 2023), which is consistent with other common hyperbolic embedding methods (Chamberlain et al., 2017; Nickel & Kiela, 2017; 2018; Sala et al., 2018), does not restrict all the features to the leaves. Assigning the features to the leaf nodes is only carried out in the tree decoding (Alg. 1), which is in line with other tree decoding algorithms (Sarkar, 2011; Chami et al., 2020; Indyk & Thaper, 2003). Note that our tree decoding algorithm generates a weighted binary tree. We assert that this does not pose a restriction because any tree can be effectively transformed into a binary tree (Bowditch, 2007; Cormen et al., 2022). For instance, consider the flexible tree in Fig. 10 (a): [sport[soccer,football,basketball]] (where “sports” is a parent node with three children: “basketball”, “soccer”, and “football”). This tree can be accurately modeled as a (weighted) binary tree as in Fig. 10 (b): [$c$[$b$[$a$[soccer, football], basketball]], sport] (where “sports”, “basketball”, “soccer”, and “football” are all leaves, $a,b,c$ are internal nodes, and the edges encode the same relationships as the flexible tree above).
@@ F.3 Differentiability of Our TWD
The proposed TWD can be expressed in a matrix form (Takezawa et al., 2021) as
@@
Note that the hyperbolic embedding in Lin et al. (2023) can be recast as a function mapping from the feature diffusion operator $\mathbf{P}$ to the product manifold of the hyperbolic spaces $\mathcal{M}$, denoted by $f_{1}:\mathbf{P}\rightarrow\mathcal{M}$. The weight $\alpha_{v}$ in the our TWD can then be expressed by
@@
F.4 The Differences between Our TWD and the Embedding Distance (Lin et al., 2023)
Our TWD is a TWD between samples that incorporates the hidden hierarchical structure of the features, while the embedding distance (Lin et al., 2023) is a hierarchical distance between the samples. Namely, in our TWD, the features are assumed to have hidden hierarchies, whereas in Lin et al. (2023), the samples are assumed to have hidden hierarchies. The work in Lin et al. (2023) does not involve TWD or OT.
More concretely, in comparison to the work in Lin et al. (2023), our work presents a new algorithm for tree decoding (Alg. 1) and a corresponding new TWD (Alg. 2). Our work also includes theoretical analysis of these two algorithms:
@@
Seemingly, from a metric learning viewpoint, both our TWD and the embedding distance (Lin et al., 2023) generate distances between the samples. Since the two are fundamentally different, the comparison can be carried out in two ways: (i) directly compute the embedding distance between the samples (denoted as ED), and (ii) use the embedding distance between the features as the ground distance for the computation of the Wasserstein distance between the samples (denoted as OT-ED). We present the classification accuracy of the word-document and scRNA-seq datasets of these two baselines in Tab. F.4. We see a clear empirical advantage of our TWD in classifying documents and scRNA-seq data compared to embedding distance (ED), as our TWD learns the latent hierarchy of the features explicitly by constructing a tree, and then, incorporating it in a meaningful distance between the samples. We note that while OT-ED performs similarly to our TWD in terms of accuracy (empirically corroborating Thm. 4.2), the computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our proposed TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. This computational advantage is empirically validated and illustrated in Fig. 4.
@@ F.5 Intuition Behind Single-Linkage Tree Construction in Hyperbolic Embeddings
The single-linkage tree aligns well with hyperbolic embeddings because it inherently emphasizes hierarchical relationships by connecting points based on the measure of similarity determined by their HD-LCAs, where there is the “treelike” nature of a hyperbolic space (Bowditch, 2007). Specifically, Prop. 4.4 shows that the LCA depth closely approximates the Gromov product, which corresponds to the tree depth in 0-hyperbolic metrics.
In the context of tree-Wasserstein distance (TWD), many tree construction methods aim to approximate a given ground metric $\widetilde{d}$ (often Euclidean) using tree metrics, enabling efficient computation of optimal transport distances. The quality of TWD methods is typically assessed by the mean relative error between the proposed TWD and the Wasserstein distance with the ground metric $\widetilde{d}$. For instance, an ultrametric incurs a $\log m$ distortion with respect to Euclidean distance (Fakcharoenphol et al., 2003), as utilized in UltraTree (Chen et al., 2024). Hence, the ultrametric Wasserstein distance (Chen et al., 2024) incurs a $\log m$ distortion compared to the original Euclidean Wasserstein distance.